OpenStudy (anonymous):

an this be solved? 96=80t-16t^2

OpenStudy (anonymous):

*can

OpenStudy (anonymous):

Yes, it can because the discriminant is not negative.

OpenStudy (anonymous):

If you put the equation into the form at^2 + bt + c = 0 which is 16t^2 - 80t + 96 = 0 you will see that if you use the quadratic solution formula:\[t = \frac{ -b \pm \sqrt{b ^{2}-4ac} }{ 2a }\]

OpenStudy (anonymous):

ooooo i completely forgot. great. thanks

OpenStudy (anonymous):

You're welcome!